[Halld-jlab] Hall D Staff Meeting Notes, October 20, 2014

Mark Ito marki at jlab.org
Mon Oct 20 21:22:55 EDT 2014


People,

Please find the notes below and at

https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_20,_2014

   -- Mark
______________________________


  Private:Hall D Staff Meeting, October 20, 2014

Present: Vladimir Berdnikov, Eugene Chudakov (chair), Mark Dalton, 
Alexandre Deur, Hovanes Egiyan, Sergey Furletov, Mark Ito, David 
Lawrence, Lubomir Pentchev, Yi Qiang, Patrizia Rossi, Elton Smith, Steve 
Smith, Sascha Somov, Simon Taylor, Beni Zihlmann


  Steve

  * section 32.10.T1 of the EH&S manual has changed
  * modifications to work planning and control
  * relevant when entering work into HDList or Atlis
  * now need detail in task hazard analysis form
  * think through step-by-step; plan appropriate hazard mitigations
  * assign risk codes, for example for
      o lock, tag, and try tasks
      o machine tools use
      o compressed gas
      o lasers
      o lead, silica, etc.
  * last page details specific changes


  Eugene

  * foreign travel: we have good insurance for needed hospital visits
  * NSAC meeting: EIC moving toward review planning, costs will be
    reviewed in Jan.


  Patrizia

  * EIC session at DNP meeting two weeks ago
      o physics w/ EIC, RHIC representative and Patrizia
      o how to build the two versions
  * PANDA: additional delay, start in 2019 or 2020.
      o funding issues
      o cost of civil construction higher than planned
      o want to start with charmonium spectroscopy


  Lubomir (DNP)

  * 4 or 5 GlueX people in attendance
  * talks from Paul Eugenio, Ryan Mitchell on XYZ spectroscopy
      o new X analysis, molecule, ccbar
  * Lambda(1405): a chi-N resonance[?]
  * new results from lattice QCD
      o resonances, form factors
      o use N-pi state for Delta
  * BNL: spin of proton
      o gluonic contribution 0.3 +- 0.1
      o orbital ang. mom. contribution small
  * Cobra[?] detector talk


  David

  * clean-up of gluon file system
      o not everything in the repository, should be
      o .tcshrc should not be changed by shift-takers
      o some use of directories not served from inside counting house
  * RAID transfer to tape library
      o IT installed software to look in a specific directory
      o we needed scripts to write there
      o auto-transfer tested, worked
      o new tape volume request to be issued to Computer Center
  * benchmarking monitoring plugins
      o CDC a memory hog
      o speed looks good
          + exception: monitoring histograms (reconstruction results)
  * can run ROOTSpy, grab macros
      o these encoded into plugins
  * running transition scripts
      o prestart, go, end
      o create directory based on run number for data
          + 20 GB per file, 300 MB/s, 1 file/minute


  Sergey F.

  * tested full chain from ROC to tape
  * working space for shift, monitoring, rate monitoring
  * calibration of flash system
  * debugging F125 firmware problems
  * Eugene: can we read out all crates? Sascha:
      o want to test trigger
      o finalize configuration files
      o BCAL, CDC crash after 1 minute at 500 Hz
      o FDC, CDC F125 fails DMA read-out
          + could exclude certain boards


  Alexandre

  * electron beam commissioning plan mainly done
      o some procedures to write
      o consistent with ACC Atlis
  * need to implement rest of commissioning in run plan
  * radiation monitors in tagger hall and Hall D in place
  * beam delayed by at least 2 days, perhaps next week
  * Yi and Alexandre will observe electron beam tuning
      o others are welcome to join
      o volunteers sign up with Y. and A.


  Mark D.

  * BCAL monitoring plugin modified, discussions with Paul Mattione and
    Kei Moriya
  * work on discriminator rates to EPICS
  * also FCAL rates from FADC
  * Eugene: what about harp scans with pair spectrometer? Hovanes:
      o can record 70 variables, CTP coincidences a candidate for EPICS
        availability


  Vladimir

  * pedestals set for BCAL, looks good
  * FDC, CDC calibrated 5 of 6 cells[?]
      o cell 4: 1 channel bad, electronics, always 0
      o fix by introducing noise by hand


  Sascha

  * BCAL trigger was working
      o cosmic trigger too
      o can select coincidence of modules
  * FCAL trigger working
      o latency screwed in some crates
      o problem receiving synch
      o will change CTP on 2 crates
      o can start and end run
          + can read-out, robust, data could be crap
  * LED FCAL trigger, cosmic self-trigger working
  * pair spectrometer trigger
      o use course PS counters, GTP, CTP
      o test with external pulser, 5 kHz with no problems
      o CTP scalers had problems, Hai Dong fixed it with firmware upgrade
  * problem with optical link PS and start counter
  * implemented start counter self-trigger
  * working on BCAL+FCAL trigger
  * format for config. files for CDC, FDC for booting decided
  * pair spectrometer ready to go, buttoned up


  Lubomir (FDC)

  * operation of F125 a concern
      o problems depending on firmware and rate
      o different versions tried
      o have list of modules reporting errors
      o will try to exclude from read-out
  * good agreement between raw and calculated data
  * pedestal calibration depends on more than one channel working; not
    always the case
  * Simon implementing histograms


  Mark

  * offline
      o simulating pure EM background with commissioning geometry
      o Kei pushing online histogram regeneration in offline jobs
  * alarm system
      o alarm limits for non-voltage systems set
      o system for version control of alarm configuration implemented
  * TOF and start counter both have self-trigger thanks to Sascha
      o configuration files a confusing area


  Simon

  * sim-recon: adding pair spectrometer hooks
  * integrating TOF event viewer from Aristeidis Tsaris
  * integrating start counter event viewer from Eric Pooser
  * need display for pair spectrometer, both tagger systems


  Elton

  * BCAL experiencing stable running
  * processed 300 people through the Hall D Safety Walk-Through!
  * Interface Control Document
      o near done
      o comments from Arne to incorporate
  * have PDUs to remotely cycle power in crates with ISEG[?] modules


  Beni

  * shares Lubomir's concern about F125s
      o cannot tell if it is hardware or firmware
      o some boards cause read-out failure, but not crash: want to reset
        these and go on
      o some boards cause crash: nothing to be done
      o spares: first article at end of month, main shipment in December
  * trigger time fed to TDCs
      o took front-panel ECL signal into ROCFDC1


  Yi

  * documentation, instructions: working with Alexandre
  * solenoid
      o received 2 spare daughter cards for power supply, will not swap
        them in
      o cryogroup wants to replace filter for cold box water
      o will not replace valve for vapor-cooled lead
  * active collimator
      o IOC for high-speed data working
      o some data loss, perhaps due to network latency, not a
        fundamental problem (timestamps exist)
      o turned the system over to UConn


  Hovanes

  * work on motorized scan GUI
      o tagger hall and Hall D
      o move, stop, count, repeat
      o same set-up as test of goniometer
      o GUIs are ready
  * beam profiler GUIs almost ready
  * finished scripts for HV alarm generation
      o tagger, FCAL missing right now
      o implemented persistency for alarm levels
  * working with Manuel on FCAL voltage control
      o Anagate devices are flaky
      o so are some bases
  * general overview GUI ready: motor positions, magnet setting,
    counting rates, etc.
  * Eugene: what about access to MCC GUIs?
      o MCC now using JTabs
      o available via ssh to MCC machine
      o Monticello also still available
      o access scheme may change in December, firewall will be erected


  Eugene

  * please send plots for upcoming reviews: event display, analysis of
    cosmic data beyond raw data
  * will have meeting to discuss how to handle ADC pedestals

Retrieved from 
"https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_20,_2014"

  * This page was last modified on 21 October 2014, at 01:19.


-- 
Mark M. Ito, Jefferson Lab, marki at jlab.org, (757)269-5295

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://mailman.jlab.org/pipermail/halld-jlab/attachments/20141020/e7b622d1/attachment.html 


More information about the Halld-jlab mailing list